OpenStudy (anonymous):

Shelly and Terrence earned points in a game by completing various tasks. Shelly completed x tasks and scored 90 points on each one. The expression below shows Terrence's total points in the game: 90x - 20 What does the constant term of the expression represent? The total points Terrence earned The total number of tasks Terrence completed The sum of Shelly's and Terrence's total points The difference between Shelly's and Terrence's total points

OpenStudy (anonymous):

pretty sure its D but i wanna make sure.

OpenStudy (anonymous):

@Night-Watcher

OpenStudy (anonymous):

''Shelly's total points would be represented by 90x. (90 points multiplied by the number of games) It is telling us that Terrence's points are 90x -20 The constant in the equation is -20. It's not A because the total is 90x-20, it's not B because we don't know how many tasks he did therefore we cannot say that it is 20, it's not the sum because not only is it a negative but it also because a total would be at least more than 90. And that leaves D as the correct answer. Because the difference between the two is 20.''

OpenStudy (anonymous):

Thank you :) I have some more :|

OpenStudy (anonymous):

Ethan is making flat rectangular pizzas for his shop. The number of slices in each pizza is 2 more than the number of pizzas he has on a single countertop. If p represents the number of pizzas on each countertop, which expression can be used to represent the number of slices on 4 countertops? 4p2 4(p + 2) 4p + 2 4p(p + 2)

OpenStudy (anonymous):

I think this one is C.

OpenStudy (anonymous):

(p+2)+(p+2)+(p+2)+(p+2) recall that a+a+a+a=4a (p+2)+(p+2)+(p+2)+(p+2)=4(p+2)

OpenStudy (anonymous):

\[4(p+2)\]

OpenStudy (anonymous):

oooh ok. Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3.

OpenStudy (anonymous):

ive got all kinds of crazy answers for this one.

OpenStudy (anonymous):

\[(-3)\times2-3\times2\times3+(-3)\times2+2\times2\times3-(-3)\times2\]

OpenStudy (anonymous):

That is the problem after all the substitutions

OpenStudy (anonymous):

kay dont tell me yet i wanna see if i can get the right answer. ill tell u what i come up with

OpenStudy (anonymous):

i got -12 but thats not any of my choices

OpenStudy (anonymous):

Whats your choices?

OpenStudy (anonymous):

omg i didnt know that it didnt make the numbers exponents im so sorry.

OpenStudy (anonymous):

Lol it's fine :)

OpenStudy (anonymous):

-a^2-3 b^3+c^2+2 b^3-c^2 where a = 3, b = 2, c = -3

OpenStudy (anonymous):

Find -a^2 - 3b^3 + c^2 + 2b^3 - c^2 if a = 3, b = 2, and c = -3.

OpenStudy (anonymous):

That's the question?

OpenStudy (anonymous):

I know the answer.

OpenStudy (anonymous):

yes thats the question.

OpenStudy (anonymous):

Choices: -67 -17 -31 1

OpenStudy (anonymous):

-17!

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

Find 3x^2 - y^3 - y^3 - z if x = 3, y = -2, and z = -5. This one has no choises.

OpenStudy (anonymous):

i got 102

OpenStudy (anonymous):

48 is your answer

OpenStudy (anonymous):

ok Simplify the expression below: a + 2c - 5b - c + a - b A.) 2a - 6b + 2c B.) 2a - 6b + c C.) a - 6b + c D.) a - 5b + 2c no exponents for this one

OpenStudy (anonymous):

2 a-6 b+c

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

I cant open that for some reason

OpenStudy (anonymous):

I cant see all the numbers

OpenStudy (anonymous):

That link says i need to log in.

OpenStudy (anonymous):

ok... work out the following 6-4(3-5)^2 + 30/5

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

How did u get that?

OpenStudy (anonymous):

I used PEMDAS to solve this problem.

OpenStudy (anonymous):

ik but what steps did you take?

OpenStudy (anonymous):

I did Parenthesis | Exponents | Multiplication | Division | Addition | Subtraction in that order.

OpenStudy (anonymous):

thats what i did but i got 28

OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. : x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

OpenStudy (anonymous):

first solve for x and then multiply the x with .6 since 60% of people downloaded Q Sam's Song.

OpenStudy (anonymous):

how do i find x?

OpenStudy (anonymous):

576

OpenStudy (anonymous):

That's your answer.

OpenStudy (anonymous):